OpenStudy (anonymous):

Integral of (2-u)/(1+u^2) du

OpenStudy (anonymous):

\[\int\limits (2-u)/(1+u^2) du\]

OpenStudy (raden):

use trig-subst it will work...

OpenStudy (raden):

do u know integration using trig-subs ?

OpenStudy (anonymous):

Not all that well. I'm looking around online for some trig identities to work and no luck so far.

OpenStudy (raden):

ok, first let u=tan(theta), so du=.... for (1+u^2) it can be equal 1+(tan(theta))^2 = ....

OpenStudy (anonymous):

I'm seeing that if u = tan(theta), then du = sec^2(theta) How does this help?

OpenStudy (raden):

yes, u are right... derivative of tan(theta) is sec^2(theta). for the second, for 1+u^2 = ... ?

OpenStudy (raden):

put u=tan(theta)

OpenStudy (anonymous):

ok... 1+(tan(theta))^2

OpenStudy (raden):

ok, from identity trigonometry : 1+(tan(theta))^2 = ... ?

OpenStudy (anonymous):

1 + tan(theta))^2 = sec^2(theta)

OpenStudy (raden):

u are correct again dont forget, for 2-u, just put u=tan(theta) so, subtitute of them to that problem , it should be ?

OpenStudy (anonymous):

\[\int\limits 2-\tan(\theta) d \theta\]

OpenStudy (raden):

yeps... just integrate that !

OpenStudy (anonymous):

\[∫2−\tan(θ)dθ = 2\theta-\ln(\sec \theta) +c\]

OpenStudy (raden):

for a moment, that's correct but the answer shoulde be in "x", right ?

OpenStudy (anonymous):

well, u really in this problem

OpenStudy (anonymous):

\[2(\arctan u) - \ln(\sec(\arctan u))+c\]

OpenStudy (raden):

ops,, sorry into "u" i mean ...

OpenStudy (raden):

from u=tan(theta), it means